OpenStudy (anonymous):

C=DA/A + 12 the fomula; A= the child age, in year D=an adult dosage C= the proper child dosage when the adult dosage is 1000 milligrams, a child is given 300 milligram. What is that child's age? Round to the nearest year.

OpenStudy (rulnick):

There may be a problem with the formula as written, particularly "DA/A"

OpenStudy (anonymous):

I'm sure I wrote it's right. which it say on the book.

OpenStudy (anonymous):

if you can't find the answer then it's fine

OpenStudy (rulnick):

The problem with "DA/A" or "D times A divided by A" is that A cancels out, and it just equals D.

OpenStudy (rulnick):

Unless you are missing parentheses: DA/(A+12)?

OpenStudy (anonymous):

if that the case then I'll try either way to find the answer.

OpenStudy (anonymous):

should this equation should look like this: C=DA/A + 12 = 300=1000/12

OpenStudy (rulnick):

In the latter case, where the denominator is A+12, you get 300 = 1000A/(A+12) 300(A+12) = 1000A 3A+36=10A 36=7A A=36/7=5 and 1/7 years

OpenStudy (rulnick):

So, to the nearest year: 5 yrs old.

OpenStudy (anonymous):

oh that the part. Thanks